# Do Teachers Dream of GenAI Widening Educational (In)equality? Envisioning the Future of K-12 GenAI Education from Global Teachers’ Perspectives

### v1

@@ 3.2. Participants: Global K–12 GenAI Education Teachers
We recruited participants based on three main criteria: (1) they were active K–12 teachers with direct classroom responsibilities; (2) they were teaching GenAI, either through dedicated AI- or IT-related courses or by incorporating GenAI into their own subject areas (e.g., English, social science, or STEM); and (3) they demonstrated a particular interest in the relationship between GenAI and educational (in)equality.
@@
[…]
@@ 4. Findings: Equality-Oriented Teaching Practices in K–12 GenAI Education (RQ1)
Figure 1. Equality-oriented GenAI teaching practices to alleviate pre-existing educational inequalities and prevent GenAI from creating new inequalities (RQ1)
@@
Across these varied approaches, teachers framed their instructions around two interconnected goals related to educational (in)equality. First, they taught students to use GenAI to mitigate existing inequalities (e.g., accessing GenAI-generated materials when textbooks or other resources were scarce) so that disadvantaged learners would not fall further behind (see subsection 4.1). Second, they aimed to prevent GenAI from creating new inequalities by teaching responsible use, critical evaluation, cultural sensitivity, and the importance of maintaining human connection (see subsection 4.2). These dual orientations show how teachers viewed GenAI as a double-edged sword. On the one hand, it could open new opportunities for students who are often excluded from advanced technologies. On the other hand, if introduced without care, it could reinforce or even widen divides (Figure 1).
@@ 4.1. Using GenAI Education to Alleviate Pre-Existing Educational Inequalities
4.1.1. Addressing Consciousness Inequality: Using GenAI Education to Bridge Gaps in Awareness and Understanding of Technology
In many schools, students had never heard of emerging technology such as GenAI or did not see why it mattered, which risked excluding them from future opportunities where AI would be essential, in the views of our sampled teachers. Teachers viewed GenAI education as a specific way to reduce this consciousness inequality by making new technology visible in students’ lives. As T6 explained, “GenAI is definitely the trend of the future… I really hope the kids can learn about it.” By introducing GenAI in her classroom, she gave students from families with little exposure to technology a first step toward recognizing its importance. Similarly, U3, a U.S. teacher nearing retirement, highlighted how teaching about GenAI pushed students to think seriously about their futures on “what are vocations like there’s certain things you really can’t replace by AI.” For both T6 and U3, teaching GenAI was about correcting unequal levels of technological awareness. Teachers used GenAI education as a direct intervention to close pre-existing gaps in students’ recognition of and engagement with emerging technology.
4.1.2. Addressing Resource Inequality: Using GenAI Education to Compensate for Scarcity of Learning Materials
@@
4.1.3. Addressing Digital Literacy Inequality: Using GenAI’s Accessibility to Broaden Computing Participation
@@
4.1.4. Personalizing and Opening Opportunities for Marginalized Students Through GenAI Education
@@ 4.2. Preventing GenAI from Creating New Inequalities Through Teaching Practice
4.2.1. New Academic Inequalities from Shortcut Use, and Teachers’ Strategies for Cultivating Critical Engagement
@@
4.2.2. Emerging Peer Inequalities from Misuse of GenAI, and Teachers’ Efforts to Set Guidance
@@
4.2.3. Cultural Biases in GenAI Outputs as a New Inequality, and Teachers’ Strategies for Critical Awareness
@@
4.2.4. Relying on GenAI for Emotional Support Instead of Interpersonal Relationships as a New Inequality, and Teachers’ Emphasis on Human Care
@@
U10 shared a particularly striking example. She explained that when she tried to manage students’ behavior, such as discouraging excessive phone use or frequent late-night outings, some students turned to ChatGPT for advice on how to negotiate with her. What surprised her even more was that students weren’t just using ChatGPT for tactical advice in dealing with teachers; they were also frequently seeking interpersonal relationship guidance and even emotional support. As a high school teacher, U10 described herself as “an at-school mom,” someone whose responsibilities extended far beyond delivering lessons. For her, teaching also meant learning to interact respectfully with others and developing the emotional resilience needed to navigate adolescence. Seeing students outsource this interpersonal and affective learning to ChatGPT made her question how the role of teachers, as moral guides, caretakers, and everyday counselors, might be shifting in the presence of GenAI tools. She began to intentionally teach students to recognize the broader social dimensions of GenAI use.
@@ 5. Findings: Continuing Inequalities Challenges in GenAI Education Beyond Individual Efforts (RQ2)
Figure 2. Continuing inequalities challenges in GenAI education beyond individual efforts (RQ2)
While section 4 highlights how teachers actively teach GenAI in their classrooms to promote equality, this section reveals the persistent barriers that limited the reach and sustainability of these teachers’ GenAI education efforts towards educational equality. These challenges were not the result of a lack of teacher commitment; indeed, many went to great depth to scaffold participation, localize tools, and bridge digital gaps. Yet, structural inequalities in AI infrastructure (see subsection 5.1), professional training (subsection 5.2), and prevailing social norms (see subsection 5.3) repeatedly constrained what teachers could achieve on their own.
@@
Many noted that these GenAI systems often reflected cultural and linguistic biases that marginalized underrepresented groups. For instance, Taiwanese teachers such as T8 and T10 found that GenAI outputs frequently defaulted to mainland Chinese vocabulary, which risked erasing local linguistic identities. South African teachers, such as those in S8, observed that GenAI tended to reproduce Western or U.S.-centric perspectives, often portraying authority in ways that excluded Black or African experiences. Additionally, most GenAI tools nowadays provide only a universal interface with little attention to accessibility. Special education teachers like T9 frequently pointed out accessibility issues such as small font size, excessive steps, and limited or poorly designed color options that prevent special education students from entering the doorway of GenAI education: “The most important thing is the font size. For special education students, larger text is better, but some GenAI tools don’t allow adjustments, so the text becomes very small.” These details are precisely what HCI researchers and GenAI designers should pay closer attention to.
@@ 6. Findings: Toward More Equal GenAI Education (RQ3)
Sampled teachers in our study reflected deeply on how they could make their own GenAI teaching more inclusive with stakeholders’ further support. These reflections ranged from advocating for school-level infrastructures that could broaden access, to suggesting company- and government-level changes that would make classroom use more accessible. While these ideas extended across multiple levels of the education ecosystem, including schools (see subsection 6.1), technology companies (see subsection 6.2), and governments (see subsection 6.3), a practical starting point was to improve individual teachers’ own GenAI teaching in the future (Figure 3).
Figure 3. Support required from other stakeholders (school, GenAI company, and the government) to better integrate GenAI education into K–12 context (RQ3)
@@ 7.1. GenAI Education as a Teacher Practice of Negotiating Educational Equality
Second, our conceptualization positions teachers as central agents of sociotechnical negotiation. While prior HCI and learning sciences literature emphasizes systemic reforms, policy initiatives, or product design as levers of educational equality (Hutson et al., 2022; Uttamchandani, 2018), our study shows that equality is also forged through the micro-practices of teachers: the activities they design, the norms they reinforce, and the cultural framings they make explicit in their classrooms. Rather than viewing teachers as passive implementers of externally designed reforms, we argue that they act as mediators of technological possibility and as guardians of educational values, deciding when GenAI functions as an equalizer and when its risks must be curtailed. This builds on critical perspectives in CSCW and HCI that foreground human actors’ interpretive work in sociotechnical systems (Xiao et al., 2025d; Cheon, 2023), extending them to the context of K–12 education by showing how teachers’ local negotiations shape the equality trajectory of AI tools.
@@ 7.2. From Traditional Computing Education to GenAI Education: New Potentials for Equalities
A second contribution of this study is to distinguish GenAI education from earlier waves of computing initiatives. Prior work on computing education often emphasized technical skill acquisition, such as programming, computational thinking, or understanding algorithmic processes (Xiao et al., 2024; Selwyn, 2019; Stamper et al., 2024; Belmar, 2022). These approaches positioned students primarily as future workers who needed to master technical competencies to remain competitive in the digital economy. By contrast, our findings show that GenAI education introduces a new paradigm: students can engage with advanced AI systems through natural language, image prompts, and conversational interaction, without requiring specialized coding knowledge or technical expertise. This shift dramatically lowers barriers to entry, enabling broader participation in AI-related learning. Teachers highlighted how GenAI’s accessibility allowed students who previously struggled with computing tasks, such as those with weak digital literacy or learning disabilities, to actively create, question, and reflect through AI-supported activities.
@@
Theoretically, this distinction builds on and extends the critical computing education literature. Orlikowski (Orlikowski, 1992) suggests that technologies are simultaneously shaped by and shaping social practices. Our findings demonstrate this duality in practice: teachers not only adopt GenAI tools but also reinterpret their role in classrooms to address equality concerns. For example, while HCI and CSCW studies (Star, 1999; Dye et al., 2019) highlight how infrastructural conditions become visible through breakdowns or exclusions, teachers in our study made GenAI’s infrastructural and cultural limitations explicit to students, turning them into teachable moments. By positioning themselves as interpreter of GenAI to their students, they shifted GenAI education into a socially situated practice of negotiating educational equality.
GenAI education, here we argued, differs from prior computing and AI literacy efforts in two key ways: (1) it opens new potentials for inclusion by lowering technical entry barriers and enabling more diverse learners to participate; and (2) it introduces new types of inequalities tied to usage practices, cultural bias, and relational dynamics that require proactive teacher mediation. These differences underscore the need to conceptualize GenAI education not as a continuation of earlier computing curricula, but as a distinct domain where the stakes of equality and social participation are more immediate and more complex.
@@ 8. Design and Policy Implications
Our findings suggest that GenAI systems must be designed with equality as a central value rather than an afterthought. Teachers repeatedly pointed out that biased outputs and inaccessible interfaces constrained their ability to provide inclusive learning experiences. This means that design cannot stop at technical efficiency or personalization but must instead address accessibility, affordability, and cultural relevance. Affordable education-oriented versions should allow meaningful experimentation and iteration rather than restricting disadvantaged learners to minimal use. Likewise, GenAI models should incorporate diverse training data and localization features that reflect local languages, cultural references, and marginalized identities, ensuring that learners in Taiwan, South Africa, or elsewhere do not feel erased by homogenizing defaults. Finally, design must move closer to pedagogy by embedding features that scaffold verification, critical reflection, and responsible classroom use, helping teachers teach not only how to operate GenAI but how to use it fairly and critically.
At the same time, teachers’ efforts were deeply circumscribed by structural conditions that cannot be resolved within the classroom alone, underscoring the need for policy interventions. Infrastructural investments are necessary to ensure that device distribution is coupled with reliable and sustainable connectivity, particularly in rural and low-income communities. Teacher capacity also requires systemic support, with professional development programs that go beyond introducing tools to offering sustained pedagogical frameworks for integrating GenAI responsibly. Moreover, our study highlights the importance of community trust: parental skepticism, cultural taboos, and political resistance shaped whether students were allowed to explore AI, suggesting that public education campaigns and participatory governance are critical for broader acceptance. Finally, equality must be pursued across institutions as well as within them, with schools encouraged to share resources, training, and expertise to prevent the Matthew Effect, where wealthy schools accelerate while others are left behind.

### v2

@@ 3.2. Participants: Global K–12 GenAI Education Teachers
We recruited participants based on three main criteria: (1) they were active K–12 teachers with direct classroom responsibilities; (2) they were teaching GenAI, either through dedicated AI- or IT-related courses or by incorporating GenAI into their own subject areas (e.g., English, social science, or STEM); and (3) they demonstrated a particular interest in the relationship between GenAI and educational (in)equality. Teachers in South Africa and the U.S. were randomly sampled through Prolific, while teachers in Taiwan were recruited via mailing lists due to the smaller eligible pool (n¡10) on Prolific. A pre-screening survey was used to ensure alignment with the three criteria.
@@
ID Location Gender Years of Experience School Type K–12 Level Teaching Subject AI-Related Course
@@ 4. Findings: Equality-Oriented Teaching Practices in K–12 GenAI Education (RQ1)
Table 2. Roles of GenAI education in alleviating existing educational inequalities and preventing the creation of new ones through teaching practices (RQ1)
GenAI Education Role Education Inequality Practices to Alleviate Inequality
Using GenAI Education to Alleviate Pre-Existing Educational Inequalities Consciousness inequality Use GenAI education to bridge gaps in awareness and understanding of technology.
Resource inequality Use GenAI education to compensate for shortages in learning materials.
Digital literacy inequality Leverage GenAI accessibility to broaden computing participation.
Lack of personalized learning for marginalized groups Open opportunities for students marginalized by traditional educational structures.
Preventing the Creation of New Inequalities Through GenAI Teaching Practices Academic inequalities from shortcut use Teacher efforts to foster critical engagement with GenAI.
Cultural biases in GenAI outputs Teacher strategies to build critical awareness of bias.
Peer inequalities from GenAI misuse Teacher efforts to set clear guidance and norms.
Excessive reliance on GenAI for emotional support Teacher efforts to emphasize human care and relationships.
@@
Across these varied approaches, teachers framed their instructions around two interconnected goals related to educational (in)equality. First, they taught students to use GenAI to mitigate existing inequalities (e.g., accessing GenAI-generated materials when textbooks or other resources were scarce) so that disadvantaged learners would not fall further behind (see subsection 4.1). Second, they aimed to prevent GenAI from creating new inequalities by teaching responsible use, critical evaluation, cultural sensitivity, and the importance of maintaining human connection (see subsection 4.2). These dual orientations show how teachers viewed GenAI as a double-edged sword. On the one hand, it could open new opportunities for students who are often excluded from advanced technologies. On the other hand, if introduced without care, it could reinforce or even widen divides (Table 2).
@@ 4.1. Using GenAI Education to Alleviate Pre-Existing Educational Inequalities
4.1.1. Addressing Consciousness Inequality: Using GenAI Education to Bridge Gaps in Awareness and Understanding of Technology.
Across all three regions, interviewees reported that many students had never encountered emerging terms such as GenAI education or AI literacy, or lacked awareness of their importance, which risked excluding them from future opportunities where AI would be essential, in the views of our sampled teachers. Teachers viewed GenAI education as a specific way to reduce this consciousness inequality by making new technology visible in students’ lives. As T6 explained, “GenAI is definitely the trend of the future… I really hope the kids can learn about it.” By introducing GenAI in her classroom, she gave students from families with little exposure to technology a first step toward recognizing its importance. Similarly, U3, a U.S. teacher nearing retirement, highlighted how teaching about GenAI pushed students to think seriously about their futures on “what are vocations like there’s certain things you really can’t replace by AI.” For both T6 and U3, teaching GenAI was about correcting unequal levels of technological awareness. Teachers used GenAI education as a direct intervention to close pre-existing gaps in students’ recognition of and engagement with emerging technology.
4.1.2. Addressing Resource Inequality: Using GenAI Education to Compensate for Scarcity of Learning Materials.
@@
4.1.3. Addressing Digital Literacy Inequality: Using GenAI’s Accessibility to Broaden Computing Participation.
@@
4.1.4. Personalizing and Opening Opportunities for Marginalized Students Through GenAI Education.
@@ 4.2. Preventing GenAI from Creating New Inequalities Through Teaching Practice
4.2.1. New Academic Inequalities from Shortcut Use, and Teachers’ Strategies for Cultivating Critical Engagement.
@@
4.2.2. Emerging Peer Inequalities from Misuse of GenAI, and Teachers’ Efforts to Set Guidance.
@@
4.2.3. Cultural Biases in GenAI Outputs as a New Inequality, and Teachers’ Strategies for Critical Awareness.
@@
4.2.4. Relying on GenAI for Emotional Support Instead of Interpersonal Relationships as a New Inequality, and Teachers’ Emphasis on Human Care.
@@
U10 shared a particularly striking example. She explained that when she tried to manage students’ behavior, such as discouraging excessive phone use or frequent late-night outings, some students turned to ChatGPT for advice on how to negotiate with her. Through conversations with students, what surprised her even more was that students were not just using ChatGPT for tactical advice in dealing with teachers; they were also frequently seeking interpersonal relationship guidance and even emotional support. As a high school teacher, U10 described herself as “an at-school mom, at-school nurses, therapists” someone whose responsibilities extended far beyond delivering lessons. For her, teaching also meant learning to interact respectfully with others and developing the emotional resilience needed to navigate adolescence. Seeing students outsource this interpersonal and affective learning to ChatGPT made her question how the role of teachers, as moral guides, caretakers, and everyday counselors, might be shifting in the presence of GenAI tools. She began to intentionally teach students to recognize the broader social dimensions of GenAI use.
@@ 5. Findings: Continuing Inequalities Challenges in GenAI Education Beyond Individual Efforts (RQ2)
Table 3. Continuing inequalities challenges in GenAI education beyond individual efforts (RQ2)
Constraint Category Description
Infrastructural Constraints on GenAI (1) Increased device availability without reliable connectivity.
Education (2) Insufficient basic digital literacy among students with limited access to technology.
(3) GenAI tool designs that introduce new forms of inequality.
Insufficient Training and Curricular (1) Uneven access to high-quality professional development for GenAI instruction.
Guidance (2) A widening gap between GenAI development and educators’ instructional preparedness.
Restrictive Social Norms (1) Resistance shaped by broader social and political norms.
(2) Traditional parental attitudes and culturally embedded beliefs.
While section 4 highlights how teachers actively teach GenAI in their classrooms to promote equality, this section reveals the persistent barriers that limited the reach and sustainability of these teachers’ GenAI education efforts towards educational equality. These challenges were not the result of a lack of teacher commitment; indeed, many went to great depth to scaffold participation, localize tools, and bridge digital gaps. Yet, structural inequalities in AI infrastructure (see subsection 5.1), professional training (see subsection 5.2), and prevailing social norms (see subsection 5.3) repeatedly constrained what teachers could achieve on their own.
@@
Many noted that these GenAI systems often reflected cultural and linguistic biases that marginalized underrepresented groups. For instance, Taiwanese teachers such as T8 and T10 found that GenAI outputs frequently defaulted to mainland Chinese vocabulary, which risked erasing local linguistic identities. South African teachers, such as S8, observed that GenAI tended to reproduce Western or U.S.-centric perspectives, often portraying authority in ways that excluded Black or African experiences: “(I asked AI to) generate a picture of a South African lady, and then, I am a Black person, it gives me a white person. I’m like, okay, make it… A black person, and now the black person looks like a thief… they don’t really have a piece of everyone in there, because they are sort of one-sided.” Additionally, most GenAI tools nowadays provide only a universal interface with little attention to accessibility. Special education teachers like T9 frequently pointed out accessibility issues such as small font size, excessive steps, and limited or poorly designed color options that prevent special education students from entering the doorway of GenAI education: “The most important thing is the font size. For special education students, larger text is better, but some GenAI tools don’t allow adjustments, so the text becomes very small.” These details are precisely what HCI researchers and GenAI designers should pay closer attention to.
@@ 6. Findings: Toward More Equal GenAI Education (RQ3)
Sampled teachers in our study reflected deeply on how they could make their own GenAI teaching more inclusive with stakeholders’ further support. These reflections ranged from advocating for school-level infrastructures that could broaden access, to suggesting company- and government-level changes that would make classroom use more accessible. While these ideas extended across multiple levels of the education ecosystem, including schools (see subsection 6.1), technology companies (see subsection 6.2), and governments (see subsection 6.3), a practical starting point was to improve individual teachers’ own GenAI teaching in the future (Table 4).
Table 4. Stakeholder-specific supports required to effectively integrate GenAI education in K–12 contexts (RQ3).
Stakeholder Support Required
School Establish Infrastructural and Organizational Support Within and Between Schools
(1) Provide equal infrastructural access to all students.
(2) Create organizational support (innovation centers and professional development) to promote teachers’ GenAI practice.
(3) Foster cross-school collaboration to redistribute cutting-edge resources among schools.
Company Create More Relatable and Accessible GenAI Education Experiences
(1) Localize GenAI products to reflect sensitive cultural topics and local norms.
(2) Design accessible GenAI education experiences.
Government Civilize GenAI Education into Society
(1) Make GenAI education a universal competency for all citizens.
(2) Co-create GenAI education with professionals across disciplines.
@@ 7.1. GenAI Education as a Teacher Practice of Negotiating Educational Equality
Second, our conceptualization positions teachers as central agents of sociotechnical negotiation. While prior HCI and learning sciences literature emphasizes systemic reforms, policy initiatives, or product design as levers of educational equality (Hutson et al., 2022; Uttamchandani, 2018), our study shows that equality is also forged through the micro-practices of teachers: the activities they design, the norms they reinforce, and the cultural framings they make explicit in their classrooms. Rather than viewing teachers as passive implementers of externally designed reforms, we argue that they act as mediators of technological possibility and as guardians of educational values, deciding when GenAI functions as an equalizer and when its risks must be curtailed. This builds on critical perspectives in CSCW and HCI that foreground human actors’ interpretive work in sociotechnical systems (Xiao et al., 2025b; Cheon, 2023), extending them to the context of K–12 education by showing how teachers’ local negotiations shape the equality trajectory of AI tools.
@@ 7.2. From Traditional Computing Education to GenAI Education: New Potentials for Equalities
A second contribution of this study is to distinguish GenAI education from earlier waves of computing initiatives. Prior work on computing education often emphasized technical skill acquisition, such as programming, computational thinking, or understanding algorithmic processes (Xiao et al., 2024; Selwyn, 2019; Stamper et al., 2024; Belmar, 2022). These approaches positioned students primarily as future workers who needed to master technical competencies to remain competitive in the digital economy. By contrast, our findings show that GenAI education introduces a new paradigm: students can engage with advanced AI systems through natural language, image prompts, and conversational interaction, without requiring specialized coding knowledge or technical expertise. This shift dramatically lowers barriers to entry, enabling broader participation in AI-related learning (Gu and Ericson, 2025). Teachers highlighted how GenAI’s accessibility allowed students who previously struggled with computing tasks, such as those with weak digital literacy or learning disabilities, to actively create, question, and reflect through AI-supported activities.
@@
Theoretically, this distinction builds on and extends the critical computing education literature. Orlikowski (Orlikowski, 1992) suggests that technologies are simultaneously shaped by and shaping social practices. Our findings demonstrate this duality in practice: teachers not only adopt GenAI tools but also reinterpret their role in classrooms to address equality concerns. For example, whereas HCI and CSCW research (Star, 1999; Dye et al., 2019) highlights how infrastructural conditions are often revealed only through breakdowns or points of exclusion, teachers in our study intentionally foregrounded GenAI’s infrastructural and cultural limits to students, transforming them into teachable moments. They positioned themselves as interpreters of GenAI, making its use a socially situated negotiation of educational equality rather than a purely technical adoption. This extends sociotechnical perspectives by showing that teachers not only respond to breakdowns, but actively curate and reinterpret GenAI’s limitations as part of equitable learning practice.
GenAI education, here we argued, differs from prior computing and AI literacy efforts in two key ways: (1) it opens new potentials for inclusion by lowering technical entry barriers and enabling more diverse learners to participate, as seen in this study and in other CHI studies work with disadvanged learners (Han et al., 2024b; Atcheson et al., 2025); and (2) it introduces new inequalities tied to usage practices, cultural bias, and relational dynamics that require proactive teacher mediation to mitigate widening disparities in different learners’ future AI readiness (Long and Magerko, 2020) and reduce exposure to AI harms (Harvey et al., 2025). These differences underscore the need to conceptualize GenAI education not as a continuation of earlier computing curricula, but as a distinct domain where the stakes of equality and social participation are more immediate and more complex.
@@ 8. Design and Policy Implications
Our findings suggest that GenAI systems must be designed with equality as a central value rather than an afterthought, as teachers repeatedly pointed out that biased outputs and inaccessible interfaces constrained their ability to provide inclusive learning experiences. Such negative experiences, on the one hand, illustrate where current systems violate key Human–AI Interaction guidelines, pointing to design issues that GenAI product designers must address (Amershi et al., 2019). More specifically, these pitfalls inform us that design cannot stop at technical efficiency or personalization but must instead address accessibility, affordability, and cultural relevance. Affordable education-oriented versions should allow meaningful experimentation and iteration rather than restricting disadvantaged learners to minimal use. Likewise, GenAI models should incorporate diverse training data and localization features that reflect local languages, cultural references, and marginalized identities, ensuring that learners in Taiwan, South Africa, or elsewhere do not feel erased by homogenizing defaults. This mirrors postcolonial HCI arguments that technologies travelling across contexts must be re-designed with local infrastructures, languages, and epistemologies in mind to avoid reinforcing global inequities (Irani et al., 2010). Finally, design must move closer to pedagogy by embedding features that scaffold verification, critical reflection, and responsible classroom use, helping teachers teach not only how to operate GenAI but how to use it fairly and critically. Moreover, teachers’ negative experiences highlight the need for stronger AI literacy among developers, managers, and decision-makers so that potential harms can be anticipated rather than treated as afterthoughts, as emphasized by CHI researchers (Xie et al., 2025).
At the same time, teachers’ efforts were deeply circumscribed by structural conditions that cannot be resolved within the classroom alone, highlighting the need for policy interventions at government, community, and school levels. Infrastructural investment is needed to ensure that device distribution is paired with reliable connectivity, particularly in rural and low-income communities-a persistent challenge highlighted in CSCW research on infrastructural inequality, especially in cultural or geographic minority contexts (Sultana et al., 2019; Star, 1999) and cannot be overlooked in future research. Teacher capacity likewise requires systemic support, with professional development that goes beyond tool introduction toward sustained pedagogical frameworks for responsible GenAI integration (Ding et al., 2024). Our findings also underscore the importance of community trust: parental skepticism, cultural taboos, and political resistance shaped students’ access to GenAI, suggesting the need for public education and participatory design, consistent with CSCW work involving community-centered design (Villari and others, 2021; Zytko et al., 2022). Finally, equity can to be addressed across schools with varied levels of resources, not only within them, according to interviewees’ success stories. Thus sharing resources and capacity-building to avoid the Matthew Effect where privileged schools accelerate while others fall behind.